OpenStudy (anonymous):

help pleaseeee write in standard form by completing the square x^2+y^2-8x+10y+15=0

OpenStudy (anonymous):

x^2-8x+y^2+10y=-15 x^2-8x+16+y^2+10y+25=-15+25+16 (x-4)^2+(y+5)^2=26